I have to find the equation for the line describes in the following in slope intercepr form
1. passing through (-4,3) (1,-7) I did this one
2. passing through (-3,-2) and paralell the thegraph of -4x+y=2 I am having problem finding the rise over run in this problem I do not know the other x and y

Answer by Alan3354(69443)

passing through (-3,-2) and paralell the thegraph of -4x+y=2 I am having problem finding the rise over run in this problem I do not know the other x and y
-----------------
For this one, find the slope, m, of the line by putting it into slope intercept form y = mx + b
-4x+y=2
y = 4x + 2
The slope is +4. All lines parallel to it will have the same slope.
Then use:
y-y1 = m*(x-x1) where (x1,y1) is the point (-3,-2)
y+2 = 4*(x+3)
y+2 = 4x+12
y = 4x + 10 slope-intercept form
4x - y = -10 standard form